# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v3

@@ 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent on the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
@@ 5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . We can infer from the modular invariance of scalar potential that the fixed points should be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega$ for the tree-level scalar potential, as the scalar potential generally diverges for $\tau\rightarrow i\infty$.
@@
It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q^{1/24}$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis. We anticipate that, due to SUSY breaking contributions (for example, from $X_{nl}$ sector), the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be altered slightly away from the fixed points $\tau_{FP}=i,\omega$. Such deviations from fixed points by SUSY breaking effects can lead to large suppression factors that regenerate $\mu_{eff}$ of order the electroweak scale.
@@
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the presence of $X_{nl}$ contributions can alter the scalar potential, making it possible to stabilize $\tau$ at an intermediate value. Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
@@
To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight $-3$ bilinear for $H_{u}$ and $H_{d}$ using proper modular forms. For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M(\tau)_{ij}$ is a modular form of level $N$ and weight $K$, the zero entries of $M(\tau)_{ij}$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point. The magnitudes of these residual symmetry breaking entries are controlled by the size of the departure $\epsilon$ from $\tau_{sym}$ and the field transformation properties under the residual symmetry group, which may depend on the modular weights Petcov:2023vws ; Novichkov:2021evw ; Petcov:2022fjf . That is, the entries of the mass matrices in the vicinity of the symmetric point are subjected to corrections of ${\cal O}(\epsilon^{l})$, where the powers $l$ corresponding to the representations of the residual symmetry group and depends only on how the representations of the fermion fields in the mass term bilinear decompose under the considered residual symmetry group. The degree of suppression, given by the integer $l$, can take values $l=0,1,\cdots,N-1$ in the case of $\tau_{sym}=i\infty$, $l=0,1,2$ in the case of $\tau_{sym}=\omega$ and $l=0,1$ in the case of $\tau_{sym}=i$. In Novichkov:2021evw , the authors systematically classify the residual symmetry representations for different choices of $\Gamma_{N}$ representations of matter fields.
@@ 5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 0$)
@@
$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}~{},$ (74)
where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (75)
$=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
@@
To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $\epsilon$. However, calculations show the suppression from $k_{3}$ power is not very efficient for obtaining a tiny number. Therefore, we choose to use a small $\epsilon$ to suppress $\mu_{eff}$ with appropriate modular weights and representations for $H_{u},H_{d}$.
@@
$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\eta(\omega)^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (77)
@@
Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and is defined by
@@
for non-cusp modular forms, where $T$ is a real constant, $D_{T}$ is the region below the line $y=T$ in the fundamental domain with $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
@@
For $\epsilon\sim 10^{-5}$, $\alpha\sim{\cal O}(1)$ and $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~{}GeV})$ in gravity mediation SUSY breaking mechanism, we can obtain a desired $\mu_{eff}$ at electroweak scale. If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, smaller $\epsilon\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$. Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be a consequence of SUSY breaking, which depends on the choice of $f_{0}$. For example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$, SUSY breaking contributions may slightly shift the minimum to some point near $\omega$. We present some benchmark points that can lead to small suppression factors in Table 5.1.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon$ $\mathcal{P}(j(\tau))$ 00 $3$ $0.2$ $0.01$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.529\times 10^{-5}$ $1$ 00 $2$ $0.2$ $0.1$ $6.758\times 10^{-6}$ $1$ 00 $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ $1$ $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.920\times 10^{-5}$ $j-4146.6830+184.0955i$ $1$ $3$ $0.2$ $0.1$ $6.587\times 10^{-5}$ $j-4144.8675+185.3946i$ Table 1: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-5}|$).
@@ 5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
[…]
To obtain the suppressed effective $\mu$-parameter, in addition to possible suppression factor from powers of $q^{1/24}$ (from $\eta(\tau)$) or $(2\Im\tau)^{-1}$ (from normalization factor), we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular functions satisfy
@@
Let us suppose that the modulus field could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account. Naively, if the desired suppression factor $10^{-20}$ to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed. Such a large $t_{0}$ possibly causes difficulties for modulus stabilization realizations, as the scalar potential for modulus field generally diverges for $\tau\rightarrow i\infty$. Therefore, it is desirable to introduce additional suppression factors other than $\epsilon$ so that medium large $t_{0}$ can lead to desired suppression factors.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (74), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (84)
@@
where proper normalization factors for $H_{u}$,$H_{d}$ from the Kahler potential are taken into account.
@@
Calculations show that the value $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.58$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale. For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}M_{SUSY}$. The tiny value of $\mu_{eff}$ results from the combined effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.

### v4

@@ 4.2 The case with $k_{X}=12$
It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent of the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
@@ 5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . As noted previously, we can infer from the modular invariance of scalar potential that the finite fixed points should in general be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega$ for the tree-level scalar potential, as the scalar potential generally diverges for $\tau\rightarrow i\infty$.
@@
It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introduce extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q^{1/24}$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
@@
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the presence of $X_{nl}$ contributions can alter the scalar potential, making it possible to stabilize $\tau$ at an intermediate value. Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
@@
To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight 3-3 bilinear for $H_{u}$ and $H_{d}$ using proper modular forms. For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M_{ij}(\tau)$ is a modular form of level $N$ and weight $K$, the zero entries of $M_{ij}(\tau)$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point. The magnitudes of these residual symmetry breaking entries are controlled by the size of the departure $\epsilon$ from $\tau_{sym}$ and the field transformation properties under the residual symmetry group, which may depend on the modular weights Petcov:2023vws ; Novichkov:2021evw ; Petcov:2022fjf . That is, the entries of the mass matrices in the vicinity of the symmetric point are subjected to corrections of ${\cal O}(\epsilon^{l})$, where the powers $l$ corresponding to the representations of the residual symmetry group and depends only on how the representations of the fermion fields in the mass term bilinear decompose under the considered residual symmetry group. The degree of suppression, given by the integer $l$, can take values $l=0,1,\cdots,N-1$ in the case of $\tau_{sym}=i\infty$, $l=0,1,2$ in the case of $\tau_{sym}=\omega$ and $l=0,1$ in the case of $\tau_{sym}=i$. In Novichkov:2021evw , the authors systematically classify the residual symmetry representations for different choices of $\Gamma_{N}$ representations of matter fields.
@@
We should note that $\mu_{eff}\ll\Lambda_{SUSY}$ can also be naturally realized through a weighton-like mechanism King:2020qaj for $H_{u}H_{d}$ bilinear. Specifically, we can introduce a single weighton-like superfield, which is a singlet under both the Standard Model and modular flavor symmetry, with a modular weight $k_{\phi}=1$. By appropriately choosing the modular weights and representations under the finite modular group $\Gamma_{N}$ for $H_{u}$ and $H_{d}$, we can ensure that only the combination
$W\supseteq\frac{Y_{\mathbf{r}}^{(k_{\mu})}}{{\left[\eta(\tau)\right]^{2k_{3}}}}\left(\frac{\phi}{M_{F}}\right)^{k}H_{u}H_{d}~,$ (104)
is allowed by modular invariance, where $k=k_{\mu}-k_{3}-k_{H_{u}}-k_{H_{d}}+3$ is a positive integer, $2k_{3}\in 24\mathbb{Z}+6$ and $M_{F}$ is a typical cutoff scale. When the weighton field $\phi$ acquires a VEV through the relevant superpotential, the electroweak-scale $\mu_{\text{eff}}$ can be naturally regenerated with a large suppression factor of order $\tilde{\phi}^{k}$, where $\tilde{\phi}\equiv\langle\phi\rangle/M_{F}\ll 1$. For example, the form of superpotential for weigthon superfield can be chosen to be King:2020qaj
$W\supseteq X(Y_{\bf 1}^{4}\frac{\phi^{4}}{M_{F}^{2}}-M^{2})~,$ (105)
with $M$ a dimensional mass parameter and $X$ an $A_{4}$ singlet. The F-flat condition $F_{X}^{*}=-W_{X}=0$ drives the weighton VEV $\langle\phi\rangle\sim(MM_{F})^{1/2}$, which leads to the suppression factor $\frac{\phi}{M_{F}}\sim(\frac{M}{M_{F}})^{1/2}\ll 1$ for $M\ll M_{F}$.
It should be noted that the $H_{u}$ and $H_{d}$ superfields, which transform non-trivially under modular symmetry, can also break the modular invariance after electroweak symmetry breaking. As the VEVs of $H_{u}$ and $H_{d}$ are much smaller than the typical scale of modulus stabilization, the extreme conditions for $\tau$ and $X_{nl}$ will receive negligible corrections from $H_{u}$ and $H_{d}$ related terms in the limits $v_{u},v_{d}\rightarrow 0$. Electroweak symmetry breaking, triggered by the VEVs of $H_{u}$ and $H_{d}$, are also affected by the modular stabilization mechanism. However, additional contributions to the electroweak symmetry breaking sector from supergravity vanish in the $M_{Pl}\rightarrow\infty$ limit, leading to negligible contributions to the low energy MSSM scalar potential for $H_{u}$ and $H_{d}$, consequently negligible shift of the VEVs for $H_{u}$ and $H_{d}$.
@@ 5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 1$)
@@
$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\left[\eta(\tau)\right]^{2k_{3}}}H_{u}H_{d}~,$ (110)
where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, $2k_{3}\in 24\mathbb{Z}+6$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\left[\eta(\omega+\epsilon)\right]^{-2k_{3}}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~,$ (111)
$=$ $\alpha\Lambda_{SUSY}\left[\eta(\omega)\right]^{-2k_{3}}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
@@
To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $|\epsilon|$. However, calculations show that the suppression from $k_{3}$ power is not very efficient to obtain a tiny number. Therefore, we choose to use a small $\epsilon$ to suppress $\mu_{eff}$ with appropriate modular weights and representations for $H_{u},H_{d}$.
@@
$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\left[\eta(\omega)\right]^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~,$ (113)
@@
Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and defined by
@@
for non-cusp modular forms, where $T$ is a real constant, $D_{T}$ is the region below the line $y=T$ in the fundamental domain with $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$. It should be noted that the normalization can not be fixed in the bottom-up approach to modular symmetry.
@@
If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, a tiny $\epsilon$ with $|\epsilon|\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$. For a smaller $\Lambda_{\text{SUSY}}$ scale, a smaller suppression factor is required to achieve an effective $\mu_{\text{eff}}$ at the electroweak scale. For instance, in gravity mediated SUSY breaking mechanism, where the SUSY breaking scale $\Lambda_{SUSY}$ is typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~GeV})$, a desired $\mu_{eff}$ at electroweak scale can be regenerated for $|\epsilon|\sim 10^{-5}$ when $\alpha\sim{\cal O}(1)$. In gauge mediated SUSY breaking mechanism where the SUSY breaking scale is as low as $10^{5}$ GeV, the required suppression factor $\epsilon$ can be of the order $|\epsilon|\sim 10^{-2}$ for $\alpha\sim\mathcal{O}(1)$.
Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be a consequence of SUSY breaking, which depends on the choice of $f_{0}$. For example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$, SUSY breaking contributions may slightly shift the minimum to some point near $\omega$.
The desired condition $|\epsilon|\ll 1$ can be achieved through typical modulus stabilization mechanisms, such as the one proposed in NPP:2201.02020 , which employs large values of $m$ and $n=0$. However, for the modulus stabilization mechanism discussed in the previous section, fine-tuning is required to obtain a minimum at $\omega+\epsilon$ with $|\epsilon|\ll 1$. This is particularly true when we consider the simplest realization ${\cal P}(j(\tau))=(j(\tau)-c_{p})^{2}$ with $c_{p}\ll 1$ for $H(\tau)=0$. We provide in Table 2 some benchmark points that can achieve small suppression factors using the aforementioned modulus stabilization mechanism. It is worth noting that the fine-tuning needed here may arise from the specific choices of input parameters for our modulus stabilization mechanism. One could explore alternative modulus stabilization mechanisms to achieve a minimum very close to $\omega$ without large fine-tuning. Moreover, such fine-tuning is not necessary when employing a weighton-type mechanism for the regeneration of $\mu_{\text{eff}}$.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon/10^{-5}$ $P(j(\tau))$
22 22 $0.2$ $0.1$ $(1.659+0.958i)$ $(j-1.765\times 10^{-9})^{2}$
0 33 $0.2$ $0.1$ $(5.441+3.141i)$ $(j-2.861\times 10^{-9})^{2}$
11 22 $0.2$ $0.1$ $(8.162+4.172i)$ $(j-3.973\times 10^{-8})^{2}$
0 22 $0.2$ $0.1$ $(6.802+3.927i)$ $(j-280.469)(j-1.008\times 10^{-8})^{2}$
22 33 $0.2$ $0.1$ $i$ $(j-1.398\times 10^{-9})^{2}$
Table 2: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon|\sim 10^{-5}|$) with $X_{nl}=0$.
@@ 5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
Although the scalar potential for the modulus generally diverges as $\tau\rightarrow i\infty$, the introduction of nilpotent Goldstino $\textbf{X}_{nl}$ superfield or multiple moduli/matter fields may alter the shape of the scalar potential at intermediate values of $\tau$, making it possible to stabilize the modulus at $\tau_{0}$ with an intermediate value of $\Im\tau_{0}$ (for example, $\Im\tau_{0}\sim 5-10$). Detailed discussions on such a possibility will be given in our subsequent studies  [^4] For example, we can choose the form ${\cal P}(j(\tau))=(j(\tau)-j(i\Im\tau_{0}))^{2}$ (or other polynomials that contain such a factor) so as that $i\Im\tau_{0}$ is a solution of ${\cal P}(j(\tau))={\cal P}^{\prime}(j(\tau))=0$, consequently being an extrema of the scalar potential. It is also possible to change the shape of scalar potential near $i\infty$ by introducing non-minimal gravitational couplings of modulus fields to Ricci scalar Nonmimimal:GR . After transforming back from the Jordan frame to Einstein frame, the modulus field can be stabilized at intermediate values of $\tau$ more easily with the transformed scalar potential.. In this work, we just concentrate on the implications of modulus stabilization at an intermediate value of $\Im\tau$ on the SUSY $\mu$-problem. Note that the stabilized VEV $\tau_{0}$ (with an intermediate value $\Im\tau_{0}$) can be regarded as lying close to the $i\infty$ fixed point.
To obtain the suppressed effective $\mu$-parameter, in addition to possible suppression factor from powers of $q^{1/24}$ (from $\eta(\tau)$) or $(2\Im\tau)^{-1}$ (from normalization factor), we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular forms satisfy
@@
Let us suppose that the modulus field could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account. Naively, if the desired suppression factor to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed. Such a large $t_{0}$ possibly causes difficulties for modulus stabilization realizations, as the scalar potential for modulus field generally diverges for $\tau\rightarrow i\infty$. Therefore, it is desirable to introduce additional suppression factors other than $\epsilon$ so that medium large $t_{0}$ can lead to desired suppression factors.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (110), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $2k_{3}\in 24\mathbb{Z}+6$, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\left[\eta(it_{0})\right]^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~,$ (120)
@@
where proper normalization factors for $H_{u}$$H_{d}$ from the Kahler potential are taken into account.
@@
Calculations show that the value $\mu_{eff}\lesssim 10^{-17}\Lambda_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.01$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale. For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}\Lambda_{SUSY}$. For a smaller $\Lambda_{SUSY}$ scale, a smaller suppression factor is needed to achieve an effective $\mu_{eff}$ at the electroweak scale, requiring a smaller value of $t_{0}$. The tiny value of $\mu_{eff}$ results from the combined effect of suppression by powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.